OpenStudy (anonymous):

What is the length of one side of a regular octagon if its perimeter is 208 mm? mm

OpenStudy (anonymous):

@Hyper

OpenStudy (hyper):

Firstly. it is known that an octogon has eight sides and the total length of these eight sides equals the perimeter which is 208mm = 20.8 cm. Therefore, length of one side = 208/8 = 20.8/8 = 26mm or 2.6 cm

OpenStudy (anonymous):

= 26mm

OpenStudy (hyper):

26mm = 2.6cm is exactly the same so it only depends on what units are required for the answer